# SweepEvGS: Event-Based 3D Gaussian Splatting for Macro and Micro Radiance Field Rendering from a Single Sweep

### v1

@@ I Introduction
Novel View Synthesis with densely multi-view images of a given 3D scene generates new and realistic 2D viewpoints [1, 2]. Recently, 3D Gaussian Splatting (3D-GS) [3] has introduced an unstructured 3D Gaussian radiance field that uses 3D Gaussian primitives to achieve impressive success in fast, high-quality, and cost-effective novel view synthesis from densely input views. However, accurate reconstruction of a radiance field typically requires capturing dense and high-quality training frames. While using videos as input can be an efficient way, blur may occur, producing poor ground truth, shown in Figure 1(a) and (c), and inaccurate camera calibration [4, 5]. Some methods [3, 6, 7, 8, 9] involve recording videos at a slow pace to minimize motion blur, ensuring that each frame is clear and usable for reconstruction. However, this approach has significant drawbacks. The process is not only slow and cumbersome but also prone to inefficiencies, particularly in dynamic environments where rapid movements are involved. Therefore, efficiently and accurately obtaining such training information has become a vital topic.
Event cameras, on the other hand, are a new imaging sensor that asynchronously captures per-pixel brightness changes, offering advanced properties and representing a major paradigm shift from conventional frame-based cameras across various vision tasks [10, 11, 12, 13, 14, 15, 16, 17]. The nature of event cameras makes them highly efficient for fast and dynamic scene capturing. The captured event streams at the microsecond level without motion blur are highly effective for dense radiance field reconstruction.
@@ II-B Neuromorphic Radiance Field Rendering
A few recent works have attempted to overcome these efficiency issues by combining 3D-GS with event data. [28] proposed a collaborative learning framework for event depth estimation, event-to-frame reconstruction, and event-based 3D-GS. However, this approach requires prior features and outputs from previous depth estimation and frame reconstruction tasks. The entire training framework also necessitates ground truth depth maps and frame references, limiting its usage, especially in scenarios where only event streams are available. [29] proposed a pure event-based 3D-GS end-to-end method for novel view synthesis. Nonetheless, the rendering quality was not satisfactory, and the method did not explore different hardware-involved imaging setups.
@@ IV-A Overview
This subsection provides an overview and structural layout of our experiments. Here, we briefly introduce the experiment setups. Experiments are conducted in three different imaging settings to demonstrate the robustness of our approach: 1) synthetic sequences, 2) real-world macro-level sequences, and 3) real-world microscopic-level sequences.
@@
For the real-world microscopic-level imaging setup, we deployed the DAVIS 346C camera on a microscope station, using the station’s mobility control to “sweep” the object horizontally or vertically. Due to manual control of the object’s movement, we could not collect paired fast-slow data as in the other setups. Therefore, we compared the visual rendering quality between the rendered result and a reference image of the object taken statically. Figure 3 shows a photograph of the setup, and detailed hardware information will be presented in Sec. IV-B.
@@
The rendering results of 3D-GS are evaluated using the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM). PSNR assesses signal fidelity by comparing the maximum signal power to noise power, while SSIM evaluates image similarity based on luminance, contrast, and structure. These metrics provide quantitative insights into the quality and fidelity of the rendered images.
@@
Figure 4: Comparison of traditional 3D-GS and our SweepEvGS on the real-world sequences under different hardware settings. Left: real-world normal sequences captured by Figure 3(a) and Figure 3(b); Right: microscopy sequences captured by Figure 3(c). For each sequence, we demonstrate (a): the event stream captured during the time interval, green for positive events, red for negative events; (b): the rendering result from 3D-GS trained on captured RGB frames; and (c): our rendering result trained on the captured event stream. Our method leverages event streams to produce sharp and clear rendering results, overcoming the blurring seen in traditional 3D-GS.
@@ IV-B Hardware Detailed Information
This subsection presents the detailed hardware information used in the real-world data collection phase to contribute to future research.
  • Event Camera: Shown in Figure 3 (a), in all real-world experiments, we use DAVIS 346 Colorful event camera, a state-of-the-art imaging device designed to capture event stream and frame data. With a frame resolution of 346 pixels by 260 pixels, this camera offers a detailed view of the environment. One of the key features of the DAVIS 346 C is it can react to events in the scene with microsecond precision. This makes it particularly effective for capturing fast-moving objects and dynamic events that would typically be blurred or missed by conventional cameras.
@@
  • Microscope: Shown in Figure 3 (c), in real-world microscopic experiments, we use Nikon Model ECLIPSE Ni-U Microscopy equipment to observe microscopic data. The Nikon Model ECLIPSE Ni-U is an advanced research-grade microscope known for its exceptional optical performance and ergonomic design. In the context of microscopy, minor shifts in the specimen or the camera can lead to considerable displacement. Such movements can render frame-based cameras ineffective, producing blurred images. Consequently, this constraint impedes conventional radiance field rendering techniques from accurately reconstructing clear and detailed textural information within microscopic environments. We manually control the movement of objects using the microscope with liner movement, making it easy to generate global coordinates. However, we did not find a way to provide precise control of the speed like the macro setup, therefore, there is no way for us to set frame correspondence between the rendered result and the ground truth frame captured in slow motion. Therefore, no quantitative comparison can be shown. However, we would love to break this hardware restriction and fill this research gap in the future.
@@ IV-C Implementation Details
Our implementation is based on the 3D-GS codebase [3], utilizing its core structure and functionalities. Since 3D-GS requires a point cloud as input, we initialize $10^{6}$ points randomly to form the initial point cloud, as the structure-from-motion initialization [4] used in the original model is not suitable for event data. All experiments were performed on an NVIDIA RTX 3090 GPU.
@@
We adhered to specific hyperparameter settings for performance and optimization. The total number of training iterations was set to 50,000. For position optimization, the initial and final learning rates were $1.6\times 10^{-4}$ and $1.6\times 10^{-6}$, respectively. The learning rates for feature, opacity, scaling, and rotation optimization were $2.5\times 10^{-3}$, $5\times 10^{-2}$, $5\times 10^{-3}$, and $1\times 10^{-3}$, respectively.
IV-D Synthetic Sequences
@@
As previously mentioned, we evaluated synthetic sequences sourced from [18], which include various scenes such as chairs, hot dogs, ficus, and mic. Table I presents a quantitative comparison with EventNeRF [27] and Ev-GS [29]  [^1] Comparison methods are evaluated with 360-degree data and the score was reported from the original paper, whereas our approach is evaluated on 90-degree data to replicate the sweep.. Our method demonstrates significant advantages over EventNeRF across these scenes. Specifically, in terms of PSNR and SSIM, as shown in Table I, our approach consistently achieves higher values, indicating superior reconstruction fidelity and better structural similarity between the reconstructed and ground truth images. Notably, our method also boasts significantly reduced training times, requiring only approximately 10 minutes compared to EventNeRF’s 14 hours for all scenes. This accelerated training time enables more efficient model development and experimentation. Moreover, our method achieves a higher temporal resolution of around 60 fps, compared to EventNeRF’s 0.32 fps. This high frame rate is crucial for smooth visual experiences, reducing motion sickness and enhancing productivity and competitiveness. Lastly, our method consumes less memory during the rendering stage, utilizing only 5 GB compared to EventNeRF’s 15 GB, which is advantageous for memory-constrained environments and facilitates scalability.
@@
TABLE I: Quantitative Comparison Against Event-NeRF [27] and Ev-GS [29] on Synthetic Sequences.
@@
IV-E Real-World Macro Sequences
TABLE II: Quantitative Comparison on Real-World Sequences. The top two results are marked with bold (1st) and underline (2nd).
Metric PSNR↑ SSIM↑ Training Time↓ FPS↑ Memory↓
Scene Toy House
EventNeRF 24.8 0.88 14h 0.30 15GB
3D-GS 23.8 0.80 7min 80.2 3GB
Ours 28.7 0.90 9min 62.7 5GB
Scene Optical Component
EventNeRF 24.1 0.81 14h 0.33 15GB
3D-GS 20.1 0.83 7min 80.9 3GB
Ours 26.3 0.87 9min 63.0 5GB
Scene Keyboard
EventNeRF 25.0 0.84 14h 0.32 15GB
3D-GS 25.3 0.86 7min 81.3 3GB
Ours 30.5 0.92 9min 62.3 5GB
Scene Cloth
EventNeRF 25.5 0.86 14h 0.32 15GB
3D-GS 24.1 0.81 7min 81.8 3GB
Ours 31.1 0.94 9min 61.3 5GB
In this experiment, we demonstrate the efficacy of our rendering method compared to others in real-world macro sequences. Due to the fast sweep during the imaging and data collection process, the captured frames are blurred. Training with blurred and low-quality frames, Traditional 3D-GS struggles, leading to blurred and indistinct renderings, as shown in part (b) in Figure 4. Our method, however, utilizes the captured event streams and the first static frame at the initial position of the camera. Event streams are, in nature, less sensitive to lighting changes. Therefore, with our designed approach, we can produce sharp and clear rendering results, as shown in sub-figure (c) in Figure 4. Our method shows a sharp rendering result with clear details and semantics. The quantitative comparison of rendering results is shown in Table II. By integrating event streams into the rendering process, our method addresses the limitations of traditional 3D-GS. The results demonstrate that our approach not only preserves but enhances the clarity and detail of the rendered images, providing a robust solution for real-world applications.
IV-F Real-World Microscopic Sequences
In microscopy settings, as demonstrated in Figure 3 part (b), even small movements of the scene or camera can cause significant displacements, resulting in blurry imaging when using frame-based cameras. This limitation often hampers traditional radiance field rendering methods from reconstructing sharp and meaningful textural information in microscopy settings. In contrast, our approach leverages event streams for radiance field rendering, making it particularly suitable for such challenging microscopy settings. As shown in the right part of Figure 4, we present two sets of visual comparisons between our approach and Frame-Based 3D-GS.
Figure 6: Visualization of the rendered results on objects from different imaging settings. The top two rows are synthetic data. The middle four rows are real-world macro objects. The bottom two rows are real-world objects captured under the microscope. We showcase the rendering results from different view perspectives for each object. We also demonstrate the rendered results on unseen views of the keyboard and cloth data. This also showcases that our approach is able to reconstruct sharp unseen views with clear semantic information.
@@ IV-G1 End-To-End Against Two Stages
To assess the necessity and effectiveness of our proposed end-to-end design, we conducted an ablation study. Specifically, we compared our rendering results to a two-stage baseline approach: first converting event data to frame data using E2VID [44] and then applying 3D-GS for rendering. Table III summarizes the quantitative evaluation results for two different scenes, Chair and Ficus, based on PSNR and SSIM metrics. As shown in Table III, our end-to-end design consistently outperforms the two-stage E2VID+3D-GS approach across both scenes and evaluation criteria. The limitations of the E2VID model become apparent when confronted with frames unseen during training. Since E2VID relies on the frames it was trained on, its intermediate results may become inaccurate when presented with novel frames. This inaccuracy propagates through the subsequent 3D-GS stage, resulting in poorer rendering quality, as illustrated in Figure 7. These results underscore the necessity and effectiveness of our end-to-end design, seamlessly integrating event-based video input processing with 3D-GS reconstruction. This integration leads to significantly improved reconstruction quality and fidelity compared to the two-stage approach.
@@
TABLE III: Ablation Study on our end-to-end approach against the two-stage baseline of using [44] plus 3D-GS. Both evaluation metrics indicates that our end-to-end approach is more effective and necessary for training and rendering across both synthesized and real-world data.

### v2

@@ I Introduction
Novel View Synthesis with densely multi-view images of a given 3D scene generates new and realistic 2D viewpoints [1, 2]. Recently, 3D Gaussian Splatting (3D-GS) [3] has introduced an unstructured 3D Gaussian radiance field that uses 3D Gaussian primitives to achieve impressive success in fast, high-quality, and cost-effective novel view synthesis from densely input views [4, 5]. However, accurate reconstruction of a radiance field typically requires capturing dense and high-quality training frames. While using videos as input can be an efficient way, blur may occur, producing poor ground truth, shown in Figure 1(a) and (c), and inaccurate camera calibration [6, 7]. Some methods [3, 8, 9, 10, 11] involve recording videos at a slow pace to minimize motion blur, ensuring that each frame is clear and usable for reconstruction. However, this approach has significant drawbacks. The process is not only slow and cumbersome but also prone to inefficiencies, particularly in dynamic environments where rapid movements are involved. Therefore, efficiently and accurately obtaining such training information has become a vital topic.
Event cameras, on the other hand, are a new imaging sensor that asynchronously captures per-pixel brightness changes [12], offering advanced properties and representing a major paradigm shift from conventional frame-based cameras across various vision tasks [13, 14, 15, 16, 17, 18, 19, 20]. The nature of event cameras makes them highly efficient for fast and dynamic scene capturing. The captured event streams at the microsecond level without motion blur are highly effective for dense radiance field reconstruction.
@@ II-B Neuromorphic Radiance Field Rendering
A few recent works have attempted to overcome these efficiency issues by combining 3D-GS with event data. [31] proposed a collaborative learning framework for event depth estimation, event-to-frame reconstruction, and event-based 3D-GS. However, this approach requires prior features and outputs from previous depth estimation and frame reconstruction tasks. The entire training framework also necessitates ground truth depth maps and frame references, limiting its usage, especially in scenarios where only event streams are available. [32, 33] proposed a pure event-based 3D-GS end-to-end method for novel view synthesis. Nonetheless, the rendering quality was not satisfactory as the work was mainly designed for synthetic event data, and the method did not explore different real-world hardware-involved imaging setups.
@@ IV-A Overview
This subsection provides an overview and structural layout of our experiments. Here, we briefly introduce the experiment setups. Experiments are conducted in three different imaging settings to demonstrate the robustness of our approach: 1) real-world macro-level sequences, 2) real-world microscopic-level sequences, and 3) synthetic sequences.
For the real-world microscopic-level imaging setup, we deployed the DAVIS 346C camera on a microscope station, using the station’s mobility control to “sweep” the object horizontally or vertically. Due to manual control of the object’s movement, we collect the event stream and RGB frames under a fast sweep and clear and sharp RGB frames under a slow sweep with a static frame capturing process to collect the RGB ground truth references. We then manually paired the fast-slow data as in the other setups. Thus, we could also compare the visual rendering quality between the rendered result and a reference image of the object taken statically under quantitative metrics. Figure 3 shows a photograph of the setup, and detailed hardware information will be presented in Sec. IV-B.
@@
The rendering results of different approaches are evaluated using the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM). PSNR assesses signal fidelity by comparing the maximum signal power to noise power, while SSIM evaluates image similarity based on luminance, contrast, and structure. These metrics provide quantitative insights into the quality and fidelity of the rendered images.
@@
Figure 4: A visual comparison of our approach against others on real-world datasets: the captured frames (a) from traditional frame-based camera results in motion blur; using traditional frame-based radiance field approach, such as 3D-GS (b) produce blurred radiance field reconstruction and view rendering; event-based NeRF approach (c), time and resource consuming while still produce accurate rendering results; further, existing event-based GS approaches (d and e) failed on real captured data as more noise and randomness are introduced in event streams; our approach (f), leveraging the proposed multi-modal data utilization and supervision, renders sharp and accurate results compared to the ground truth references (g). We visually compare the results across all methods on real-world macro sequences (toy, optical component, keyboard, and cloth) captured by hardware in Figure 3(a) and Figure 3(b); and also on real-world microscopy sequences captured by hardware in Figure 3(c). Dotted boxes indicate areas of interest, and the solid line boxes reveal zoomed results.
@@ IV-B Hardware Detailed Information
This subsection presents the detailed hardware information used in the real-world data collection phase to contribute to future research. A video is included as supplementary material to demonstrate the hardware setups used in this paper.
  • Event Camera: Shown in Figure 3 (a), in all real-world experiments, we use DAVIS 346 Colorful event camera, a state-of-the-art imaging device designed to capture event stream and frame data. With a frame resolution of 346 pixels by 260 pixels, this camera offers a detailed view of the environment. One of the key features of the DAVIS 346 C is that it can react to events in the scene with microsecond precision. This makes it particularly effective for capturing fast-moving objects and dynamic events that would typically be blurred or missed by conventional cameras.
@@
  • Microscope: Shown in Figure 3 (c), in real-world microscopic experiments, we use Nikon Model ECLIPSE Ni-U Microscopy equipment to observe microscopic data. The Nikon Model ECLIPSE Ni-U is an advanced research-grade microscope known for its exceptional optical performance and ergonomic design. In the context of microscopy, minor shifts in the specimen or the camera can lead to considerable displacement. Such movements can render frame-based cameras ineffective, producing blurred images. Consequently, this constraint impedes conventional radiance field rendering techniques from accurately reconstructing clear and detailed textural information within microscopic environments. We manually control the movement of objects using the microscope with linear movement, making it easy to generate global coordinates. However, we did not find a way to provide precise control of the speed like the macro setup, therefore, we manually aligned the fast sweep frame and the slow sweep ground truth frame to compute quantitative metrics.
@@ IV-C Implementation Details
Our implementation is based on the 3D-GS codebase [3], utilizing its core structure and functionalities. The project was built using Python and CUDA toolkit [48]. Since 3D-GS requires a point cloud as input, we initialize $10^{6}$ points randomly to form the initial point cloud, as the structure-from-motion initialization [6] used in the original model is not suitable for event data. All experiments were performed on an NVIDIA RTX 3090 GPU.
@@
We adhered to specific hyperparameter settings for performance and optimization. The total number of training iterations was set to 50,000. For position optimization, the initial and final learning rates were $1.6\times 10^{-4}$ and $1.6\times 10^{-6}$, respectively. The learning rates for feature, opacity, scaling, and rotation optimization were $2.5\times 10^{-3}$, $5\times 10^{-2}$, $5\times 10^{-3}$, and $1\times 10^{-3}$, respectively. Densification parameters included a percent density of $0.01$, a densification interval of $100$ iterations, and densification occurring between iterations $500$ and $50,000$, triggered by a gradient threshold of $2\times 10^{-4}$. The opacity was reset every $3,000$ iterations. These settings were chosen from the default 3D-GS settings to balance training stability and performance.
IV-D Real-World Macro Sequences
TABLE I: Quantitative Comparison on Real-World Sequences. The top two results are marked with bold (1st) and underline (2nd).
Metric PSNR↑ SSIM↑ Training Time↓ FPS↑ Memory↓
Scene Toy House
3D-GS 23.8 0.80 7min 80.2 3GB
EventNeRF 24.8 0.88 14h 0.30 15GB
Ev-GS 19.2 0.68 10min 59.7 5GB
Event-3DGS 26.7 0.85 15min 55.2 7GB
Ours 28.7 0.90 9min 62.7 5GB
Scene Optical Component
3D-GS 20.1 0.83 7min 80.9 3GB
EventNeRF 24.1 0.81 14h 0.33 15GB
Ev-GS 20.2 0.70 10min 55.3 5GB
Event-3DGS 25.8 0.82 15min 53.0 7GB
Ours 26.3 0.87 9min 63.0 5GB
Scene Keyboard
3D-GS 25.3 0.86 7min 81.3 3GB
EventNeRF 25.0 0.84 14h 0.32 15GB
Ev-GS 13.6 0.60 10min 61.2 5GB
Event-3DGS 25.2 0.79 15min 57.0 7GB
Ours 30.5 0.92 9min 62.3 5GB
Scene Cloth
3D-GS 24.1 0.81 7min 81.8 3GB
EventNeRF 25.5 0.86 14h 0.32 15GB
Ev-GS 11.6 0.58 10min 55.9 5GB
Event-3DGS 26.4 0.77 15min 53.2 7GB
Ours 31.1 0.94 9min 61.3 5GB
In this experiment, we demonstrate the efficacy of our rendering method compared to others in real-world macro sequences. Due to the fast sweep during the imaging and data collection process, the captured frames are blurred. Training with blurred and low-quality frames, Traditional 3D-GS struggles, leading to blurred and indistinct renderings, as shown in part (b) in Figure 4. Existing event-based approaches such as EventNeRF [30], Ev-GS [32], and Event-3DGS [33], on the other hand, are mainly designed for synthetic event data. Therefore, when dealing with real-world event streams that are noisy and with more randomness, they failed to produce high-quality radiance fields and rendering results, shown in parts (c, d, and e) in Figure 4. Our method, however, utilizes the captured event streams and the first static frame at the initial position of the camera. Event streams are, in nature, less sensitive to lighting changes. Therefore, with our designed approach, we can produce sharp and clear rendering results, as shown in sub-figure (f) in Figure 4. Our method shows a sharp rendering result with clear details and semantics. The quantitative comparison of rendering results is shown in Table I. Our method addresses the limitations of traditional 3D-GS and challenges faced by existing event-based radiance field approaches. The results demonstrate that our approach not only preserves but enhances the clarity and detail of the rendered images, providing a robust solution for real-world applications.
IV-E Real-World Microscopic Sequences
In microscopy settings, as demonstrated in Figure 3 part (b), even small movements of the scene or camera can cause significant displacements, resulting in blurry imaging when using frame-based cameras. This limitation often hampers traditional radiance field rendering methods from reconstructing sharp and meaningful textural information in microscopy settings. In contrast, our approach leverages event streams for radiance field rendering, making it particularly suitable for such challenging microscopy settings. The quantitative comparison results are shown in Table II, and our approach renders high-quality reconstruction results with decent PSNR, SSIM scores, while keeping a balanced efficiency in training and rendering. As shown in the right part of Figure 4, we present visual comparisons between our approach, Frame-Based 3D-GS, and existing event-based radiance field approaches.
TABLE II: Quantitative Comparison on Real-World Microscopy Sequences. The top two results are marked with bold (1st) and underline (2nd).
Metric PSNR↑ SSIM↑ Training Time↓ FPS↑ Memory↓
Scene Onion
3D-GS 23.8 0.80 7min 80.2 3GB
EventNeRF 24.8 0.88 14h 0.30 15GB
Ev-GS 8.6 0.55 10min 62.0 5GB
Event-3DGS 18.9 0.75 15min 57.0 7GB
Ours 28.7 0.90 9min 62.7 5GB
Scene Fly
3D-GS 20.1 0.83 7min 80.9 3GB
EventNeRF 24.1 0.81 14h 0.33 15GB
Ev-GS 9.8 0.55 10min 60.0 5GB
Event-3DGS 24.1 0.82 15min 55.0 7GB
Ours 26.3 0.87 9min 63.0 5GB
Figure 5: More visualization of the rendered results on objects from different imaging settings. The top two rows are synthetic data. The middle four rows are real-world macro objects. The bottom two rows are real-world objects captured under the microscope. We showcase the rendering results from different view perspectives for each object. We also demonstrate the rendered results on unseen views of the keyboard and cloth data. This also showcases that our approach is able to reconstruct sharp unseen views with clear semantic information.
IV-F Synthetic Sequences
@@
As previously mentioned, we evaluated synthetic sequences sourced from [21], which include various scenes such as chairs, hot dogs, ficus, and mic. Table III presents a quantitative comparison with EventNeRF [30] and Ev-GS [32]  [^1] Comparison methods are evaluated with 360-degree data and the score was reported from the original paper, whereas our approach is evaluated on 90-degree data to replicate the sweep.. Our method demonstrates significant advantages over EventNeRF across these scenes. Specifically, in terms of PSNR and SSIM, as shown in Table III, our approach consistently achieves higher values, indicating superior reconstruction fidelity and better structural similarity between the reconstructed and ground truth images. Notably, our method also boasts significantly reduced training times, requiring only approximately 10 minutes compared to EventNeRF’s 14 hours for all scenes. This accelerated training time enables more efficient model development and experimentation. Moreover, our method achieves a higher temporal resolution of around 60 fps, compared to EventNeRF’s 0.32 fps. This high frame rate is crucial for smooth visual experiences, reducing motion sickness and enhancing productivity and competitiveness. Lastly, our method consumes less memory during the rendering stage, utilizing only 5 GB compared to EventNeRF’s 15 GB, which is advantageous for memory-constrained environments and facilitates scalability.
@@
TABLE III: Quantitative Comparison Against Event-NeRF [30] and Ev-GS [32] on Synthetic Sequences.
@@ IV-G1 End-To-End Against Two Stages
To assess the necessity and effectiveness of our proposed end-to-end design, we conducted an ablation study. Specifically, we compared our rendering results to a two-stage baseline approach: first converting event data to frame data using E2VID [49] and then applying 3D-GS for rendering. Table IV summarizes the quantitative evaluation results for two different scenes, Chair and Ficus, based on PSNR and SSIM metrics. As shown in Table IV, our end-to-end design consistently outperforms the two-stage E2VID+3D-GS approach across both scenes and evaluation criteria. The limitations of the E2VID model become apparent when confronted with frames unseen during training. Since E2VID relies on the frames it was trained on, its intermediate results may become inaccurate when presented with novel frames. This inaccuracy propagates through the subsequent 3D-GS stage, resulting in poorer rendering quality, as illustrated in Figure 7. These results underscore the necessity and effectiveness of our end-to-end design, seamlessly integrating event-based video input processing with 3D-GS reconstruction. This integration leads to significantly improved reconstruction quality and fidelity compared to the two-stage approach.
@@
TABLE IV: Ablation Study on our end-to-end approach against the two-stage baseline of using [49] plus 3D-GS. Both evaluation metrics indicates that our end-to-end approach is more effective and necessary for training and rendering across both synthesized and real-world data.